OpenStudy (anonymous):

Simplify. –(2x – 3y) + 5x + 2 – 3x + 3y + 2 A. 4 B. 4x + 4 C. 6y + 4 D. 4x + 6y + 4

OpenStudy (shamil98):

Distribute -1 to 2x and -3y after that combine all the like terms to get your answer.

OpenStudy (anonymous):

D ?

OpenStudy (shamil98):

-2x + 3y + 5x + 2 -3x +3 y + 2 -2x + 5x -3x = 0 3y + 3y = 6y 2 + 2 = 4 6y + 4

OpenStudy (anonymous):

oh i messed up at the -2x